# Continual Learning for Encoder-only Language Models via a Discrete Key-Value Bottleneck

### v1

@@ 5 Experimental Setup
In our continual learning experiments, we compare the performance of DKVB to other CL methods in the three settings described above, namely TIL, DIL, and CIL. In addition, we adapt the challenging single-head CIL setup from Träuble et al. (2023) to topical text classification. We take the best-performing bottleneck architectures from the pre-experiments and apply the same bottleneck parameters and hyperparameters. We use accuracy as the primary evaluation metric in all our experiments and present the average performance and standard deviation over five runs. These runs involve random initialization and randomized task sequence order. Additionally, we report the average per epoch runtime of each method. Details about the implementation, hyperparameters, and bottleneck parameters can be found in the Appendix A.2.
@@
For the main experiments, we use three datasets, two of which has been also used in Ke et al. (2021). In the DIL setting, we use the Document Sentiment Classification (DSC) dataset. It consists of $10$ subsets of product reviews with a positive or negative sentiment label. Each subset constitutes a separate task with $4,000$ training, $500$ validation, and $500$ test samples. Since the tasks are similar and only differ in the product domain, this dataset is used to evaluate knowledge transfer. In the CIL setup, we use the earlier described 20ng dataset Lang (1995). Similarly to Ke et al. (2021), we create a sequence of $10$ tasks consisting of two classes each. This setup is mainly used to test the models ability to overcome catastrophic forgetting. For the TIL setup, we create a sequence of tasks by combining 4 tasks from the GLUE benchmark Wang et al. (2018). This dataset which we call 4GLUE includes four different task-types: The RTE dataset is used for testing natural language inference, the MRPC is used for measuring semantic textual similarity, the SST-2 is a popular dataset for sentiment analysis, and the QQP dataset which is used for natural language inference and question answering. For the single-head CIL experiments, we use two different versions of the R21578 news dataset Lewis (1997), R8 (includes 8 classes) and R52 (with 52 classes). Due to the R21578 dataset having a highly skewed class frequency distribution, we simulate a low-resource training scenario and include only $100$ samples from each class in both datasets. On R8 we divide the dataset into 8 increments with one class in each increment. On R52 we create 26 increments, each with two random classes.
@@
DKVB-NP Incr. BERT (frozen) $80.99_{\textit{2.07}}$ $79.58_{\textit{1.91}}$ $59.67_{\textit{1.59}}$ $54.58_{\textit{1.66}}$ $58.12_{\textit{2.89}}$ $50.14_{\textit{1.90}}$
@@
DKVB-P Incr. BERT (frozen) $74.09_{\textit{4.88}}$ $68.01_{\textit{5.10}}$ $57.81_{\textit{2.00}}$ $52.89_{\textit{2.77}}$ $58.77_{\textit{3.23}}$ $51.02_{\textit{1.81}}$
@@
We follow the CL evaluation method by De Lange et al. (2021). The model is trained sequentially on all tasks and is evaluated by averaging out the test performance of each task recorded after the final training increment. This results in each task in the sequence being a binary classification problem. In the multi-head configurations (for CIL and TIL), we use a separate decoder for each task and provide the task-ID during training and evaluation. To further investigate the continual learning capabilities of DKVB, we implement the single-head CIL setup of Träuble et al. (2023). Compared to the multi-head CIL task, this setup is considered to be more challenging and lacks explicit task boundaries. For its evaluation, the models are tested on the whole test data after each increment, including previously unseen classes.
@@
We use the best-performing methods reported in Ke et al. (2021), selecting one from each CL approach (cf. Section 2.1). From regularization-based methods, we choose EWC (Serra et al., 2018), a common baseline with strong performance in many CL studies. DER++ (Buzzega et al., 2020) belongs to the replay-based methods and uses distilled knowledge from past experiences to guide the incremental training process. OWM (Zeng et al., 2019) is an optimization-based approach that constrains the gradient updates to a direction orthogonal to the input space of previously trained tasks. Lastly, CTR (Ke et al., 2021) is an architecture-based approach, that utilizes capsule networks to prevent catastrophic forgetting and facilitate knowledge transfer. We also include three baselines without any additional forgetting or knowledge transfer handling, noted as naive continual learning (NCL).
For DKVB we take the best-performing architectures from Section 3.3, and include both the parametric (DKVB-P) and non-parametric (DKVB-NP) variants. We experiment with three different strategies for key initialization. In the first two strategies we use the training data for initializing the keys: in the incremental setup the keys are optimized in a continual fashion before each task using only the training data of the given increment (denoted as Incr.), while in the full initialization setup the keys are initialized once before training, using the full training input distribution (denoted as Oracle). In the third setup, we use a cross-domain corpus different from the training data, with the goal of creating keys that are general purpose (denoted as Generic). For this, we use a small version of the English Wikipedia dump [^1] https://huggingface.co/datasets/wikipedia, which is commonly included in pre-training datasets. In all three setups, we use an EMA decay of 0.2. For the Incr. and Oracle setups the key initialization is set to three epochs, while for the Generic we use 1 epoch.
All CL methods (except CTR) are applied to a frozen BERT model and have a single-head configuration without any task-ID information for the DIL scenario and a multi-head configuration with task-ID provision on the CIL and TIL scenarios. CTR is based on an Adapter-BERT (Houlsby et al., 2019) backbone and requires a multi-head setup and task-ID information for its dynamic architecture in all scenarios. For the single-head CIL experiments, we include the naive baselines and the replay-based DER++ method. The rest of the CL baseline methods either require explicit task boundaries for optimal performance (OWM, EWC) or only work in a multi-head configuration (CTR).
@@ 6 Results
The results of the main experiments are shown in Table 2. In the DIL setting the difference in accuracy between the baseline methods is low, with CTR having the highest score of $88.73\%$. The performance of the DKVB variants in this scenario is below the baselines. In the CIL setting, there is a substantial variation between model performance, with half of the CL methods achieving over 90% accuracy, while BERT NCL, DER++, and the incremental DKVB variants having an accuracy score below $60\%$. The best result on the CIL dataset was achieved with the non-parametric DKVB Oracle ($97.06\%$) followed by EWC ($96.80\%$) and BERT frozen NCL ($96.35$%). Within the TIL scenario, the highest accuracy scores were achieved with BERT frozen NCL ($71.90\%$) and CTR ($72.71\%$). Within the DKVB variants the best performance was consistently seen with the non-parametric Oracle variant, closely followed by the non-parametric Generic variant. On the CIL and TIL scenarios both of these methods outperformed most of the baselines. Additional measures of the backward transfer performances can be found in Appendix A.4.
@@
To compare the computational costs of the different methods we measure the average epoch runtimes for each model. The results can be found in Table 3. The DKVB without key initialization is the closest in speed to the NCL with a frozen BERT where only a parametric decoder is optimized during training. While the regularization-based EWC and the optimization-based OWM methods also achieve comparable runtime to the NCL frozen BERT model, the addition of DER++ or CTR results in a substantial increase in runtime. The key initialization process scales with the number of samples used (hence it takes longer in the Oracle and Generic setups), but overall the computational requirements for DKVB with key initialization is still lower than most other CL methods as the Oracle and Generic key initializations have to be run only once.
Table 3: Epoch runtimes and standard deviations (in subscript) in seconds averaged over a single run.
CL Method Model DIL (DSC) CIL (20NG) TIL (4GLUE) NCL BERT $20.6_{\textit{3.0}}$ $8.9_{\textit{0.0}}$ $482.3_{\textit{659.4}}$ NCL BERT (frozen) $4.4_{\textit{0.6}}$ $1.9_{\textit{0.0}}$ $105.5_{\textit{144.2}}$ NCL Adapter-BERT $24.1_{\textit{3.4}}$ $10.4_{\textit{0.0}}$ $566.2_{\textit{772.7}}$ DER++ BERT (frozen) $26.2_{\textit{0.9}}$ $7.4_{\textit{2.5}}$ $249.7_{\textit{361.2}}$ EWC BERT (frozen) $8.0_{\textit{0.08}}$ $2.3_{\textit{0.2}}$ $129.0_{\textit{176.6}}$ OWM BERT (frozen) $6.7_{\textit{0.3}}$ $2.0_{\textit{0.1}}$ $108.9_{\textit{148.6}}$ CTR Adapter-BERT $487.1_{\textit{0.4}}$ $195.2_{\textit{0.1}}$ $3011.7_{\textit{0.0}}$ DKVB-NP BERT (frozen) $4.67_{\textit{0.6}}$ $2.00_{\textit{0.0}}$ $109.35_{\textit{149.2}}$ DKVB-P BERT (frozen) $4.88_{\textit{0.7}}$ $2.07_{\textit{0.0}}$ $114.38_{\textit{156.5}}$ Key Initialization in DKVB Incremental $4.7_{\textit{0.6}}$ $1.9_{\textit{0.0}}$ $111.6_{\textit{152.5}}$ Oracle $46.9_{\textit{0.5}}$ $19.5_{\textit{0.1}}$ $446.0_{\textit{0.9}}$ Generic $469.0_{\textit{2.1}}$ $469.0_{\textit{2.1}}$ $469.0_{\textit{2.1}}$
@@
The results of the single-head class incremental learning are shown in Figure 2. The highest accuracy score is 81.17% on R8 and 47.78% on R52. Both scores were achieved with the non-parametric DKVB variant using the Generic and the Oracle key initialization, respectively. On both datasets, the non-DKVB models, which included the BERT frozen NCL and DER++ displayed sharp drops in performance between increments, indicating the occurrence of catastrophic forgetting and overfitting on current training increment. DER++ showcased better performance than the naive baseline, but still underperforms the Oracle and Generic variants, with a final accuracy score of 16.70% on R8 and 35.75% on R52. The detailed results with additional models (BERT NCL, EWC) can be found in Appendix A.4.
@@ 7 Discussion
Our experiments show that introducing a discrete key-value bottleneck benefits CL with encoder-only language models, both in terms of performance and efficiency. A non-parametric decoder is consistently preferable for CL. The best key initialization is obtained by unsupervised access to the full input feature distribution, but utilizing a general-purpose corpus for key initialization is also a viable option for NLP tasks. Below, we discuss these key insights.
@@
We found that employing pooling before the bottleneck has a substantial negative effect on the model performance (see Section 3.3). This suggests that in contrast to lower dimensional vision tasks Träuble et al. (2023), it is necessary to retain the full dimensionality of the text encodings. Similarly, CLS pooling is inferior compared to mean pooling across all setups. Segmenting on the token dimension only worked in the case of parametric decoding, indicating that the output of the DKVB module yields better representational power if the segmentation happens on the hidden dimension. An additional linear layer after the DKVB module as a parametric decoder can compensate for encodings with weaker representational power. However, a non-parametric decoder produces comparable results in most configurations.
@@
In the CL experiments reported in Section 5, the non-parametric DKVB variants achieved comparable results to the other CL methods with runtimes on par with the NCL frozen BERT variant. In the CL scenarios, incremental key initialization consistently led to subpar performance. This indicates that access to a general purpose corpus or to the full input distribution is needed to achieve competitive performance with DKVB. The largest performance drop between the DKVB variants and other CL methods was seen in the DIL setting. This suggests that DKVB is less suitable for knowledge transfer, and is more adept at alleviating catastrophic forgetting. Additionally, we showed that NCL methods achieve similar results as the CL methods in the DIL setup. This indicates that pre-trained language models are already adept at the task of domain incremental learning. In the CIL and TIL tasks, only the frozen BERT NCL variant showcased comparable performance to the CL methods. The strong performance of frozen BERT in these experiments suggests that if task-ID is available during testing, a simple multi-head configuration with a frozen encoder is often sufficient. Experiments in the single-head CIL setup have shown to be more challenging. As the models are tested on the full test set after each increment, ideally they should exhibit a progressive increase in accuracy. But when no task-ID is provided and decoding is done with a single head, most models overfit and suffer catastrophic forgetting between increments. The only model that showcased improved CL capability in this scenario was the DKVB. While having access to the full data distribution may be unrealistic in practice, our experiments show that this is not needed in NLP. Rather, when initializing the keys using a general-purpose corpus, we obtain results close to the Oracle setup.
@@ 8 Conclusion
The discrete key-value bottleneck (DKVB) offers a unique and efficient approach for continual learning (CL). It enables context-dependent updates in the model without explicit parameter isolation or by dynamically expanding the architecture. Considering the special challenges of CL with text embeddings, we analyzed twelve architectural variants of the bottleneck. The best variants apply mean pooling after the bottleneck and utilize the hidden dimension of the encoded input representation for creating the bottleneck heads. We conducted a comprehensive evaluation across different continual learning settings in NLP, i.e., domain-incremental learning, class-incremental learning, and task-type incremental learning, and showed that DKVB with proper key initialization offers consistent improvement in most settings and is comparable to dedicated CL methods from the literature. Moreover, we showed that DKVB can be used even in the most challenging single-head continual learning scenarios when no task-ID is provided.

### v2

@@ 5 Experimental Setup
In our continual learning experiments, we compare the performance of DKVB to other CL methods in the three settings described above, namely TIL, DIL, and CIL. In addition, we adapt the challenging single-head CIL setup from Trauble et al. Träuble et al. (2023) to topical text classification. We take the best-performing bottleneck architectures from the pre-experiments and apply the same bottleneck parameters and hyperparameters. We use accuracy as the primary evaluation metric in all our experiments and present the average performance and standard deviation over five runs. These runs involve random initialization and randomized task sequence order. Additionally, we report the average per epoch runtime of each method. Details about the implementation, hyperparameters, and bottleneck parameters can be found in the Appendix B.
@@
For the main experiments, we use three datasets, two of which have also been used in Ke et al. (2021). We use the Document Sentiment Classification (DSC) dataset in the DIL setting. It consists of 101010 subsets of product reviews with a positive or negative sentiment label. Each subset constitutes a separate task with 4,0004,0004 , 000 training, 500500500 validation, and 500500500 test samples. Since the tasks are similar and only differ in the product domain, this dataset is used to evaluate knowledge transfer. In the CIL setup, we use the earlier described 20ng dataset Lang (1995). Similarly to Ke et al. (2021), we create a sequence of 101010 tasks consisting of two classes each. This setup is mainly used to test the models’ abilities to overcome catastrophic forgetting. For the TIL setup, we create a sequence of tasks by combining four tasks from the GLUE benchmark Wang et al. (2018). This dataset, which we call 4GLUE, includes four different task-types: The RTE dataset is used for testing natural language inference, the MRPC is used for measuring semantic textual similarity, the SST-2 is a popular dataset for sentiment analysis, and the QQP dataset which is used for natural language inference and question answering.
For the single-head CIL experiments, we use two different versions of the R21578 news dataset Lewis (1997), R8 (includes 8 classes) and R52 (with 52 classes). Due to the R21578 dataset’s highly skewed class frequency distribution, we simulate a low-resource training scenario and include only 100100100 samples from each class in both datasets. On R8, we divide the dataset into 8 increments, with one class for each increment. On R52, we create 26 increments, each with two random classes.
@@
DKVB-NP Incremental BERT (frozen) $80.99_{\textit{2.07}}$ $79.58_{\textit{1.91}}$ $59.67_{\textit{1.59}}$ $54.58_{\textit{1.66}}$ $58.12_{\textit{2.89}}$ $50.14_{\textit{1.90}}$
@@
DKVB-P Incremental BERT (frozen) $74.09_{\textit{4.88}}$ $68.01_{\textit{5.10}}$ $57.81_{\textit{2.00}}$ $52.89_{\textit{2.77}}$ $58.77_{\textit{3.23}}$ $51.02_{\textit{1.81}}$
@@
We follow the CL evaluation procedure of De Lange et al. (2021). A model is trained sequentially on all tasks and is evaluated by averaging the test performance of each task recorded after the final training increment. This results in each task in the sequence being a binary classification problem. In the multi-head configurations (for CIL and TIL), we use a separate decoder for each task and provide the task ID during training and evaluation. To further investigate the continual learning capabilities of the DKVB, we implement the single-head CIL setup of Träuble et al. (2023). Compared to the multi-head CIL task, this setup is considered to be more challenging and lacks explicit task boundaries. For its evaluation, the models are tested on the whole test data after each increment, including previously unseen classes.
@@
We use the best-performing methods reported in Ke et al. (2021), selecting one from each CL approach (cf. Section 2.1). From regularization-based methods, we choose EWC Serra et al. (2018), a common baseline with strong performance in many CL studies. DER++ Buzzega et al. (2020) belongs to the replay-based methods and uses distilled knowledge from past experiences to guide the incremental training process. OWM Zeng et al. (2019) is an optimization-based approach that constrains the gradient updates to a direction orthogonal to the input space of previously trained tasks. Lastly, CTR Ke et al. (2021) is an architecture-based approach that utilizes capsule networks to prevent catastrophic forgetting and facilitate knowledge transfer. We also include three baselines without any additional forgetting or knowledge transfer handling, noted as naive continual learning (NCL).
For DKVB we take the best-performing architectures from Section 3.3, and include both the parametric (DKVB-P) and non-parametric (DKVB-NP) variants. We experiment with three different strategies for key initialization. In the first two strategies, we use the training data for initializing the keys: in the incremental setup, the keys are optimized in a continual fashion before each task using only the training data of the given increment (denoted as Incremental), while in the full initialization setup, the keys are initialized once before training, using the full training input distribution (denoted as Oracle). In the third setup, we use a cross-domain corpus different from the training data to create general-purpose keys (denoted as Generic). For this, we use a small version of the English Wikipedia dump [^2] https://huggingface.co/datasets/wikipedia, which is commonly included in pre-training datasets. In all three setups, we use an EMA decay of 0.20.20.2. For the Incremental and Oracle setups, the key initialization is set to three epochs, while for the Generic we use one epoch.
All CL methods (except CTR) are applied to a frozen BERT model and have a single-head configuration without any task-ID information for the DIL scenario and a multi-head configuration with task-ID provision on the CIL and TIL scenarios. CTR is based on an Adapter-BERT Houlsby et al. (2019) backbone and requires a multi-head setup and task-ID information for its dynamic architecture in all scenarios. For the single-head CIL experiments, we include the naive baselines and the replay-based DER++ method. The rest of the CL baseline methods either require explicit task boundaries for optimal performance (OWM, EWC) or only work in a multi-head configuration (CTR).
@@ 6 Results
[…]
@@
We measure the average epoch runtimes for each model to compare the computational costs of the different methods. The results can be found in Table 3. Among the evaluated methods, DKVB achieves runtime closest to NCL with a frozen BERT, where training is limited to optimizing a parametric decoder. While the regularization-based EWC and the optimization-based OWM methods also achieve a runtime comparable to the NCL frozen BERT model, adding replay in DER++ and dynamic architecture in CTR substantially increases runtime. The key initialization process scales with the number of samples, but the overall computational cost of DKVB remains lower than most continual learning methods since initialization is done once before training and involves just a forward pass. The average runtime of key initialization is shown in Table 4
Table 3: Per-epoch training runtimes (in seconds), averaged over a single run. Standard deviations are shown as subscripts.
CL Method Model DIL (DSC) CIL (20NG) TIL (4GLUE) NCL BERT $20.6_{\textit{3.0}}$ $8.9_{\textit{0.0}}$ $482.3_{\textit{659.4}}$ NCL BERT (frozen) $4.4_{\textit{0.6}}$ $1.9_{\textit{0.0}}$ $105.5_{\textit{144.2}}$ NCL Adapter-BERT $24.1_{\textit{3.4}}$ $10.4_{\textit{0.0}}$ $566.2_{\textit{772.7}}$ DER++ BERT (frozen) $26.2_{\textit{0.9}}$ $7.4_{\textit{2.5}}$ $249.7_{\textit{361.2}}$ EWC BERT (frozen) $8.0_{\textit{0.08}}$ $2.3_{\textit{0.2}}$ $129.0_{\textit{176.6}}$ OWM BERT (frozen) $6.7_{\textit{0.3}}$ $2.0_{\textit{0.1}}$ $108.9_{\textit{148.6}}$ CTR Adapter-BERT $487.1_{\textit{0.4}}$ $195.2_{\textit{0.1}}$ $3011.7_{\textit{0.0}}$ DKVB-NP BERT (frozen) $4.67_{\textit{0.6}}$ $2.00_{\textit{0.0}}$ $109.35_{\textit{149.2}}$ DKVB-P BERT (frozen) $4.88_{\textit{0.7}}$ $2.07_{\textit{0.0}}$ $114.38_{\textit{156.5}}$
Table 4: Per-epoch key initialization runtimes (in seconds) and corresponding sample sizes. Standard deviations are shown as subscripts.
Key Initialization DIL (DSC) CIL (20NG) TIL (4GLUE) Incremental $4.7_{\textit{0.6}}$ $1.9_{\textit{0.0}}$ $111.6_{\textit{152.5}}$ (n=4 000) (n=1 600) (n=87 470) Oracle $46.9_{\textit{0.5}}$ $19.5_{\textit{0.1}}$ $535.22_{\textit{1.9}}$ (n=40 000) (n=16 000) (n=349 881) Generic $469.0_{\textit{2.1}}$ $469.0_{\textit{2.1}}$ $469.0_{\textit{2.1}}$ (n=205 328) (n=205 328) (n=205 328)
@@
The single-head class incremental learning results are shown in Figure 2. The highest accuracy scores are 81.17% on R8 and 47.78% on R52. Both scores were achieved with the non-parametric DKVB variant using the Generic and Oracle key initialization, respectively. On both datasets, the non-DKVB models, which included the BERT frozen NCL and DER++, displayed sharp drops in performance between increments, indicating the occurrence of catastrophic forgetting and overfitting on the current training increment. DER++ showcased better performance than the naive baseline but still underperformed the Oracle and Generic variants, with a final accuracy score of 16.70% on R8 and 35.75% on R52. The detailed results with additional models (BERT NCL, EWC) can be found in Appendix Section C.
@@ 7 Discussion
Our experiments show that fine-tuning encoder-only language models with an optimal discrete key-value bottleneck architecture achieves comparable results to partial fine-tuning in standard learning scenarios, but greatly benefits CL, both in terms of performance and efficiency. The best key initialization is obtained by unsupervised access to the full input feature distribution, but utilizing a general-purpose corpus for key initialization is also a viable option for NLP tasks. Below, we discuss these key insights.
@@
We found that employing pooling before the bottleneck has a substantial negative effect on the model performance (see Section 3.3). This suggests that in contrast to lower dimensional vision tasks Träuble et al. (2023), it is necessary to retain the full dimensionality of the text encodings. Similarly, CLS pooling is inferior compared to mean pooling across all setups. Segmenting on the token dimension only worked in the case of parametric decoding, indicating that the DKVB module’s output yields better representational power if the segmentation happens on the hidden dimension. An additional linear layer after the DKVB module, acting as a parametric decoder, can compensate for encodings with weaker representational power. However, a non-parametric decoder produces comparable results in most configurations.
@@
In the CL experiments reported in Section 5, the non-parametric DKVB variants achieved comparable results to other CL methods and maintained runtimes on par with the NCL frozen BERT variant, but only when using pre-initialized keys.
When using incremental key initialization, performance was consistently subpar, indicating that DKVB requires access to a general-purpose corpus or the full input distribution to achieve competitive performance. While having access to the full data distribution may be unrealistic in practice, our experiments show that this is not needed in NLP. Rather, when initializing the keys using a general-purpose corpus, we obtain results that are close to the Oracle setup.
The largest performance drop between the DKVB variants and other CL methods was seen in the DIL setting. This suggests that DKVB’s strength in preventing catastrophic forgetting through distinct key-value bindings becomes its weakness in DIL, as this compartmentalization restricts the model’s ability to transfer knowledge across different domains. Notably, NCL methods achieve similar results as the CL methods in this DIL setup, indicating that pre-trained language models without a bottleneck are already well-suited for domain incremental learning.
In the CIL and TIL tasks, only the frozen BERT NCL variant showcased performance comparable to that of the CL methods. The strong performance of frozen BERT in these experiments suggests that if task-ID is available during testing, a simple multi-head configuration with a frozen encoder is often sufficient. Experiments in the single-head CIL setup have shown to be more challenging. As the models are tested on the full test set after each increment, ideally, they should exhibit a progressive increase in accuracy. But when no task-ID is provided and decoding is done with a single head, most models overfit and suffer catastrophic forgetting between increments, with DKVB being the only model to demonstrate improved CL capability in this scenario. This suggests that DKVB’s unique architecture effectively maintains knowledge across tasks without needing task-specific heads.
@@ 8 Conclusion
The discrete key-value bottleneck offers an efficient approach to continual learning. It enables context-dependent updates in the model without explicit parameter isolation or dynamically expanding the architecture. Considering the special challenges of continual learning with text embeddings, we analyzed twelve architectural variants of the bottleneck. The best variants apply mean pooling after the bottleneck and utilize the hidden dimension of the encoded input representation to create the bottleneck heads.
We conducted a comprehensive evaluation across different continual learning settings in NLP, i. e., domain-incremental learning, class-incremental learning, and task-type incremental learning, and showed that with proper key initialization, the discrete key-value bottleneck offers consistent improvement in most settings and is comparable to dedicated continual learning methods from the literature. Moreover, we showed that it can be used even in the most challenging single-head continual learning scenarios when no task-ID is provided.